tickets to a concert cost either 12 or 15 dollars a total 300 tickets are sold and the total receipts were 4140 dollars how many of each kind of tickets were sold
Answer by sachi(548)

let x tickets of $12 & (300-x) tickets of & 15 were sold for a total value of $4140
so 12x+15(300-x)=4140
or 12x+4500-15x=4140
or 3x=4500-4140=360
or x=360/3=120
so 120 tickets of $12 & 180 tickets of & 15 were sold
ans